OpenStudy (anonymous):

50 less than three times a number is equal to the number. What is the number?

OpenStudy (anonymous):

3x-50=x x=25

OpenStudy (amistre64):

3n - 50 = n 2n = 50 n = 50/2

OpenStudy (anonymous):

25